Pd=20e^0.4Q and Ps=100e^-0.4Q find the equilibrium price and quantity
1
Expert's answer
2021-07-30T04:07:01-0400

According to microeconomics at equilibrium

"Ps = Pd"

"20e^{0.4Q}= 100e^{-0.4Q}"

"20e^{0.8Q}= 100"

"e^{0.8Q} = \frac{100}{20}"

"0.8Q = ln 5"

"Q= \frac{5}{4} ln 5"

Use the value of Q in the equation


"20^{0.4(5\/4 ln5)}"

"20^{0.5 ln 5}"

Equilibrium price = "20e^{0.5ln 5}"

Quantity= "\frac{5}{4} ln 5"